OpenStudy (anonymous):

Nina divided 36 by a number and the result was between 0 to 3. Which integer could have been the divisor?

OpenStudy (anonymous):

-6 18 6 -18 chooses

OpenStudy (anonymous):

Divide 36 by 18=2

OpenStudy (anonymous):

Try dividing 36 by 0, 1, 2, 3. 36/0 doesn't make sense 36/1 = 36 36/2 = 18 36/3 = 12 36, 18 and 12 are all potential divisors.

OpenStudy (anonymous):

This works because for any number: n / a = b The following also works: n / b = a